SpaceRef

SpaceRef


NOAA SATOPS Morning Report: Thursday, January 12, 2012

Status Report From: NOAA Office of Satellite Operations
Posted: Friday, January 13, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:

- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 012

GOES
GOES-12 DOY011 frame breaks at 14:02z and 21:03z.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety

ESPC Reported data outage and delays
GOES-13 (East) 1145z and 1215z images to be missed. ESPC has switched to the backup NPORT.

Beginning January 11, 2012 @ 0015z ESPC has not been receiving HRPT data from Monterey. Monterey (FNMOC) is investigating the problem.

DMSP SSMI/S data delayed; specifically DMSP F-16, F-17 and F-18 EDRR, SDRR, TDRR and TDUP. As of 2255z ESPC began receiving all backlogged DMSP SSMI/S data.

Questions or comments to Ron.Mahmot@noaa.gov

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks